Parallel rays of green mercury light with a wavelength of 546 nm pass through a slit with a width of 0.437 mm. What is the distance from the central maximum to the first minimum on a screen 1.75 m away from the slit?
Answer #1

λ = 546 nm, w = 0.437 mm, L = 1.75 m, find the distance d

d = λL/w = 2.18*10-3 m = 2.18 mm